### v3

# A GREAT Architecture for Edge-Based Graph Problems Like TSP

Abstract
In the last years, an increasing number of learning-based approaches have been proposed to tackle combinatorial optimization problems such as routing problems. Many of these approaches are based on graph neural networks (GNNs) or related transformers, operating on the Euclidean coordinates representing the routing problems. However, such models are ill-suited for a wide range of real-world problems that feature non-Euclidean and asymmetric edge costs. To overcome this limitation, we propose a novel GNN-based and edge-focused neural model called Graph Edge Attention Network (GREAT). Using GREAT as an encoder to capture the properties of a routing problem instance, we build a reinforcement learning framework which we apply to both Euclidean and non-Euclidean variants of vehicle routing problems such as Traveling Salesman Problem, Capacitated Vehicle Routing Problem and Orienteering Problem. Our framework is among the first to tackle non-Euclidean variants of these problems and achieves competitive results among learning-based benchmarks.
I Introduction
Graph neural networks (GNNs) have emerged as a powerful tool for learning on graph-structured data such as molecules, social networks, or citation graphs [1]. In recent years, GNNs have also been applied in the setting of combinatorial optimization, especially routing problems [2, 3, 4] since such problems can be modeled as graph problems. Alternatively, transformer models [5] have been used as encoder architectures for routing problems in learning-based settings [6, 7, 8]. Both architectures, GNNs and transformers, are used to iteratively compute node embeddings of the problem instances. They do so through message-passing operations and the attention mechanism, respectively, starting from the Euclidean coordinates and, optionally, other node features like customer demands as initial features. However, as a result, these architectures are limited to Euclidean problem instances.
In reality, routing problems are often non-Euclidean or even asymmetric due to, e.g., traffic congestion (and time being the objective), one way streets (and distance being the objective) or elevation changes (and energy consumption being the objective). Non-Euclidean routing problems are specified by pair-wise distances between the nodes (e.g., in form of distance matrices). Such pair-wise distances are examples of graph edge features and can not be represented as node features like Euclidean coordinates. As a result, since GNNs and transformers focus on operating on and computing node-level features, they are not well-suited for such asymmetric problems.
In this paper, we overcome the limitations of regular GNNs and transformers by introducing the Graph Edge Attention Network (GREAT). This results in the following contributions:
  • Whereas traditional GNNs are node-focused because of their node-level message passing operations, GREAT is purely edge-focused, meaning information is passed along edges sharing endpoints. This makes GREAT perfect for edge-level tasks such as routing problems. We note, however, that the idea of GREAT is task-independent and can potentially also be applied in other suitable edge-focused settings, possibly chemistry or road networks.
  • We utilize GREAT in a reinforcement learning (RL) framework that can be trained end-to-end to predict optimal tours of the Traveling Salesman Problem (TSP), Capacitated Vehicle Routing Problem (CVRP) and Orienteering Problem (OP). As the inputs of GREAT are edge features (e.g., distances), GREAT applies to all variants of such routing problems, including non-Euclidean variants like the asymmetric TSP. The resulting trained framework achieves state-of-the-art performance for two different asymmetric distribution in TSP, CVRP and OP with 100 nodes. Notably, our approach outperforms MatNet [9] which is the current standard for non-Euclidean problems.
  • We propose a few-shot curriculum learning (CL) framework to fine-tune our trained GREAT models from TSP with 100 nodes to bigger instance sizes. It is based on incrementally increasing the problem instance size up to 500 nodes and, by this, familiarizing the obtained models with larger TSP instances.
  • In an additional experimental evaluation, we investigate the generalization performance of GREAT. We show that GREAT generalizes well to larger TSP instances (even in a zero-shot setting without CL) and achieves competitive performance for asymmetric TSP up to $1,000$ nodes. Further, we show that GREAT is also applicable to real-world routing problems, achieving strong generalization performance on TSPLIB, CVRPLIB and OPLIB.
The remainder of this article is structured as follows: we introduce relevant background and literature in Section II. Afterwards, GREAT is introduced in Section III. Our GREAT-based RL framework, the exact experimental setup and its results are presented in Section IV before we conclude in Section V.
II Preliminaries
II-A Graph Neural Networks
Graph neural networks are a class of neural architectures that operate on graph-structured data. In contrast to classical neural network architectures, where the neuron connections are fixed and grid-shaped, the connections in a GNN reflect the structure of the input data.
GNNs iteratively compute node feature vectors by aggregating over the node feature vectors of adjacent nodes and mapping the old feature vector together with the aggregation to a new node feature vector. Additionally, the feature vectors are multiplied with trainable weight matrices and nonlinearities are applied to achieve actual learning. The node feature vectors of a neighborhood are typically scaled in some way (depending on the respective GNN architecture) and, sometimes, edge feature vectors are also considered within the aggregations. A conceptual example can be found in Figure 1. For the mathematical formulation of such an iterative node feature update, we consider the Graph Attention Network (GATs [10]) whose update formula is given by:
[…]
where the attention score $\alpha_{i,j}$ between node $i$ and $j$ that determines the scaling (or “importance”) of a neighboring node is computed by
[…]
While the original GAT does not incorporate edge features in its message passing operations, some extensions of the architecture integrate both node and edge features for improved representation learning. For example, [11] propose Edge-Featured Graph Attention Networks (EGAT) which uses edge features by applying a GAT not only on the input graph itself but also its line graph representation (compare [12] as well) and then combining the computed features. Another work incorporating both node and edge features in the node-level aggregations in an attention-based GNN is [13] who use a “Graph Transformer” for a semi-supervised classification task. Further, [14] introduce “EdgeFormers”, an architecture operating on Textual-Edge Networks where they combine the success of Transformers in LLM tasks and GNNs. Their architecture also augments GNNs to utilize edge (text) features. By this, text representations (e.g., the text of a product review corresponding to edge $(i,u)$ of an item $i$ by a user $u$) can be enriched by structural graph data. We note, however, that none of these architectures are purely edge-feature-focused.
[…]
Figure 1: Classical GNN: node feature is updated by aggregating over the neighboring nodes’ feature vectors and, optionally, adjacent edges
II-B Routing Problems
II-B1 Traveling Salesman Problem
The traveling salesman problem (TSP) is an NP-hard combinatorial optimization problem. Given a set of nodes (representing “cities”) and the pairwise distances between them, he objective of the problem is to find a shortest route that visits all nodes exactly once and returns to the starting node in the end. We provide an example of an Euclidean TSP (where the nodes are distributed in the Euclidean plane) including a possible solution tour in Figure 2(a). The depicted tour was generated using the well-known LKH heuristic [15]. In practice, TSP instances are often asymmetric. Therefore, in this paper, we consider not only Euclidean TSP instances (EUC; with coordinates sampled uniformly at random in the unit square) but additionally TSP instances of the TMAT distribution (following [9]). TMAT leads to asymmetric distances that follow the triangle inequality, which means that for three nodes $u,v,w$ in our set of nodes it holds that $d(u,v)\leq d(u,w)+d(w,v)$. We follow the TMAT distribution generation of [9] but normalize the distance matrices such that the biggest distance between two nodes is exactly 1. Further, we consider the XASY distribution (extremely asymmetric) where all pairwise distances are sampled uniformly at random from the range $(0,1)$ which means that the triangle inequality between different nodes does generally not hold, making the problem more challenging. The same distribution has been used before in [16].
II-B2 Capacitated Vehicle Routing Problem
In the capacitated vehicle routing problem (CVRP), we are given a set of customer nodes with associated demands. Further, we have a special depot node and a vehicle with a certain, fixed capacity. The goal of the problem is to find a route that starts and ends in the depot, visits all customers exactly once to deliver “goods” in accordance to their demands and minimizes the traveled distance. Since in general the total amount of customer demands exceeds the vehicle capacity, it is usually necessary to intermediately return to the depot several times. An example of a Euclidean CVRP instance with a possible tour can be found in Figure 2(b). The tour was generated using the well-known HGS heuristic [17]. In our work, we consider CVRP instances where the pairwise distances between nodes are sampled from the same EUC, TMAT and XASY distributions used for TSP. Further, we use the customer demands and vehicle capacity used in [6], which means for 100 customers we have a vehicle with a capacity of 50 and the demands are sampled uniformly at random from $\{1,\dots,9\}$.
II-B3 Orienteering Problem
In the orienteering problem (OP), each node (representing a customer) has a collectible prize. Further, we have a vehicle that starts in a designated depot location. The goal of the problem is to find a tour that maximizes the prices collected from the the customers while not traveling further than a specified maximum distance (and, again, it is not allowed to visit a customer more than once). In the end, the vehicle returns to the depot node. We provide an example of a Euclidean OP instance with 50 customers in Figure 2(c). The solution to the problem was found using the EA4OP heuristic [18]. In this work, the distances between the nodes in our OP instances follow the earlier introduced EUC, TMAT and XASY distributions. The prizes that can be collected by visiting the customers are sampled using a similar uniform specification to [6], i.e., we assign prizes sampled uniformly from $\{\frac{1}{100},\dots,\frac{100}{100}\}$ to our customers. The maximum travel distance (considering instances with 100 customers) is set to $4$ for EUC and TMAT instances and to $0.4$ for XASY instances.
(a) TSP instance with 50 nodes
(b) CVRP instance with 50 customers
(c) OP instance with 50 customers
Figure 2: Routing problem examples
II-C Learning to Route
In recent years, many studies have tried to solve routing problems such as the TSP or CVRP with learning-based methods.
Popular approaches for solving routing problems with the help of machine learning include RL frameworks, where encodings for the problem instances are computed. These encodings are then used to incrementally build solutions by selecting one node in the problem instance at a time. Successful works in this category include [19, 20, 6, 7, 8].
Another possibility to use machine learning for solving routing problems is to predict edge probabilities or scores which are later used in search algorithms such as beam search or guided local search. Examples for such works are [2, 6, 21, 4, 3, 22, 23].
A further possibility involves iterative methods where a solution to a routing problem is improved over and over until a stopping criterion (e.g., convergence) is met. Possibilities for such improvements are optimizing subproblems or applying improvement operators such as $k$-opt. Examples for such works are [24, 25, 26, 27, 28, 29].
II-C1 Non-Euclidean Routing Problems
Many of the mentioned works, especially in the first two categories, use GNNs or transformer models (which are related to GNNs via GATs [30]) to capture the structure of the routing problem in their neural architecture. This is done by interpreting the coordinates of Euclidean routing problem instances as node features. These node features are then processed in the GNN or transformer architectures to produce encodings of the overall problem. However, this limits the applicability of such works to Euclidean routing problems. This is unfortunate, as non-Euclidean routing problems are highly relevant in reality [31].
So far, only a few studies have also investigated non-Euclidean versions of routing problems, such as the asymmetric TSP (ATSP). One such study is [16] where they solve synthetic ATSP instances with unsupervised learning, RL, and supervised learning approaches. Another study is [32] that uses online RL to solve ATSP instances of TSPLIB [33]. Another successful work tackling ATSP is [9]. There, the Matrix Encoding Network (MatNet) is proposed, a neural model suitable to operate on matrix encodings representing combinatorial optimization problems such as the distance matrices of (A)TSP. Their model is trained using RL. [34] proposes a modular framework to solve larger routing problem instances and uses MatNet as a subsolver to tackle bigger (A)TSP instances. [35, 36] proposes a learning formulation that allows training on relatively small routing problem instances while still performing well on larger instances at inference. Further, they propose a model that is trained for several routing problems at once. While doing so, they also tackle ATSP, using a “normal” GNN as their encoder that uses random one-hot vectors as node-feature inputs together with edge features that encode the actual distances.
III Graph Edge Attention Network
[…]
Figure 3: GREAT (node-free): edge “attends” to adjacent edges
Existing GNNs are based on node-level message-passing operations. In contrast, we propose an edge-level-based GNN where information is compared between and passed along neighboring edges. This makes our model tailored for edge-level tasks such as edge classification (e.g., in the context of routing problems, determining if edges are “promising” to be part of the optimal solution or not). Our model is attention-related, meaning the “focus” of an edge to another adjacent edge in the update operation is determined using an attention-inspired mechanism. Consequently, similar to the Graph Attention Network (GAT) we call our architecture Graph Edge Attention Network (GREAT). A simple visualization of the idea of GREAT is shown in Figure 3. In this visualization, edge $e_{14}$ compares information between (or “attends” to) all other edges it shares an endpoint with. While GREAT is a task-independent framework, it is suited perfectly for routing problems: Consider (A)TSP as an example. In this setting, node features are typically absent, only edge features are given as distances between nodes. A standard node-level GNN would not be well-suited to process such data. Existing work uses coordinates of the nodes in the Euclidean space as node features to overcome this limitation (e.g., [2, 6]). However, this strategy is limited to Euclidean TSP and does not extend to other forms of the problem, like asymmetric cases. GREAT, however, can be applied to all TSP variants.
While in theory it would also be possible to construct a line graph to transform edges into nodes and by this make “normal” GNNs applicable to the problem, this would transform an original dense (A)TSP graph with $n$ cities and $n^{2}$ edges into a representation with $\mathcal{O}(n^{2})$ nodes and $\mathcal{O}(n^{3})$ edges (compare [3]), increasing the order of magnitude of the representation by one.
We note that even though GREAT has been developed in the context of routing problems, it generally is a task-independent architecture and it might be useful in completely different domains as well such as chemistry, road, or flow networks which fall beyond the scope of this study.
III-A Architecture
In the following, we provide the mathematical model defining the different layers of a GREAT model. In particular, we propose two variants of GREAT.
The first variant is purely edge-focused and does not have any node features. Here, each edge compares information with all other edges it shares at least one endpoint with. The idea essentially corresponds to the visualization in Figure 3. In the following, we refer to this variant as “node-free” GREAT or GREAT “NF”.
The second variant is also edge-focused but has intermediate, temporary node features. This essentially means that nodes are used to save all information on adjacent edges. Afterwards, the features of an edge are computed by combining the temporary node features of their respective endpoints. These node features are then deleted and not passed on to the next layer, only the edge features are passed on. The idea of this GREAT variant is visualized in Figures 4 and 5. In the remainder of this study, we refer to this GREAT version as “node-based” or GREAT “NB”.
[…]
Figure 4: GREAT (node-based): compute temporary node features
[…]
Figure 5: GREAT (node-based): combine temporary node features
GREAT NB vs. GREAT NF
The advantages and disadvantages of both GREAT variants focus mostly on technical implementation aspects. While both GREAT variants were implemented in Python using Pytorch [37], the GREAT NB implementation requires additional message-passing operations by Pytorch Geometric [38] which makes the architecture slightly slower. However, the GREAT NB implementation generalizes more easily to non-complete, sparse graphs. Moreover, we hypothesize that the NB variant ensures a more considerable information exchange between adjacent edges compared to the purely edge-based variant. We note that the code (as well as the trained models and datasets of the experiments in Section IV) are publicly available. [^1] https://github.com/attila-lischka/GREAT
III-B Mathematical Formulations
We now describe the mathematical equations defining the internal operations of GREAT. Similar to the original transformer architecture of [5], GREAT consists of two types of sublayers: attention sublayers and feedforward sublayers. We always alternate between attention and feedforward sublayers. The attention sublayers can be node-based (with temporary nodes features) or completely node-free. Using the respective sublayers leads to overall node-based or node-free GREAT. A visualization of the block stacking can be found in Figure 6.
GREAT NB, Attention Sublayers: For each node in the graph, we compute a temporary node feature in layer $\ell$ as
$x_{i}^{\ell}=\sum_{j\in N(i)}(\alpha^{\prime}_{i,j}W^{\prime}_{1}e_{i,j}^{\ell-1}||\alpha^{\prime\prime}_{i,j}W^{\prime\prime}_{1}e_{j,i}^{\ell-1})$
where
$\alpha^{\prime}_{i,j}$ $=\text{softmax}\big(\frac{(W^{\prime}_{2}e_{i,j}^{\ell-1})^{\top}W^{\prime}_{3}e_{i,j}^{\ell-1}}{\sqrt{d}}\big)$
$=\frac{\exp\big(\frac{(W^{\prime}_{2}e_{i,j}^{\ell-1})^{\top}W^{\prime}_{3}e_{i,j}^{\ell-1}}{\sqrt{d}}\big)}{\sum_{k\in N(i)}\exp\big(\frac{(W^{\prime}_{2}e_{i,k}^{\ell-1})^{\top}W^{\prime}_{3}e_{i,k}^{\ell-1}}{\sqrt{d}}\big)}$
$\alpha^{\prime\prime}_{i,j}$ $=\text{softmax}\big(\frac{(W^{\prime\prime}_{2}e_{j,i}^{\ell-1})^{\top}W^{\prime\prime}_{3}e_{j,i}^{\ell-1}}{\sqrt{d}}\big)$
$=\frac{\exp\big(\frac{(W^{\prime\prime}_{2}e_{j,i}^{\ell-1})^{\top}W^{\prime\prime}_{3}e_{j,i}^{\ell-1}}{\sqrt{d}}\big)}{\sum_{k\in N(i)}\exp\big(\frac{(W^{\prime\prime}_{2}e_{k,i}^{\ell-1})^{\top}W^{\prime\prime}_{3}e_{k,i}^{\ell-1}}{\sqrt{d}}\big)}$
Note that we compute two attention scores and concatenate the resulting values to form the temporary node feature. This allows GREAT to differentiate between incoming and outgoing edges which, e.g. in the case of asymmetric TSP, can have different values. If symmetric graphs are processed (where $e^{\ell}_{i,j}=e^{\ell}_{j,i}$ for all nodes $i,j$) we can simplify the expression to only one attention score.
The temporary node features are concatenated and mapped to the hidden dimension again to compute the actual edge features of the layer.
$e^{\ell}_{i,j}=W_{4}(x^{\ell}_{i}||x^{\ell}_{j})$
We note that $W^{\prime}_{1},W^{\prime\prime}_{1},W^{\prime}_{2},W^{\prime\prime}_{2},W^{\prime}_{3},W^{\prime\prime}_{3},W^{\prime}_{4},W^{\prime\prime}_{4}$ are trainable weight matrices of suitable dimension, $d$ is the hidden dimension and $||$ denotes concatenation. $W^{\prime}_{1}e_{i,j}^{\ell-1},W^{\prime}_{2}e_{i,j}^{\ell-1}$ and $W^{\prime}_{3}e_{i,j}^{\ell-1}$ are similar to the “values”, “keys” and “queries” of the original transformer architecture.
GREAT NF, Attention Sublayers: Here, edge features of network layer $\ell$ are directly computed as
$e_{i,j}^{\ell}=(\alpha^{\prime}_{i,j}W^{\prime}_{1}e_{i,j}^{\scriptscriptstyle\ell-1}||\alpha^{\prime}_{j,i}W^{\prime}_{1}e_{j,i}^{\scriptscriptstyle\ell-1}||\alpha^{\prime\prime}_{i,j}W^{\prime\prime}_{1}e_{i,j}^{\scriptscriptstyle\ell-1}||\alpha^{\prime\prime}_{j,i}W^{\prime\prime}_{1}e_{j,i}^{\scriptscriptstyle\ell-1})$
Note that the edge feature consists of four individual terms that are concatenated. Due to the attention mechanism, these terms summarize information on all edges outgoing from node $i$, ingoing to node $i$, outgoing from node $j$, and ingoing to node $j$. The distinction between in- and outgoing edges is again necessary due to asymmetric graphs. The $\alpha^{\prime}$ and $\alpha^{\prime\prime}$ scores are computed as for the node-based GREAT variant above.
Remark that the attention used in GREAT (NB and NF) could also be interpreted as a form of gating and it is more similar to the type of “attention” used in architectures like the attention free transformer proposed by [39] that do not use self-attention. Computing real self-attention scores (as used in the original transformer paper [5]) would require us to compare each of the $\mathcal{O}(n^{2})$ edges in a routing problem graph to its $\mathcal{O}(n)$ adjacent edges. This would result in calculating $\mathcal{O}(n^{3})$ many attention scores (compared to $\mathcal{O}(n^{2})$ the way we do it) and lead to a considerable computational increase.
Feedforward (FF) Sublayer: Like in the original transformer architecture, the FF layer has the following form.
$e^{\ell}_{i,j}=W_{2}\text{ReLU}(W_{1}e_{i,j}^{\ell-1}+b_{1})+b_{2}$
where $W_{1},b_{1},W_{2},b_{2}$ are trainable weight matrices and biases of suitable sizes. Moreover, again like in [5], the feedforward sublayers have internal up-projections, which temporarily double the hidden dimension before scaling it down to the original size.
We further note that we add residual layers and normalizations to each sublayer (attention and FF). Therefore, the output of each sublayer is (like in the original transformer architecture):
$e^{\ell}_{i,j}=\text{LayerNorm}(e_{i,j}^{\ell-1}+\text{Sublayer}(e_{i,j}^{\ell-1}))$
Edge Feature Inputs GREAT Sublayer Add & Norm Feed Forward Add & Norm Edge EmbeddingsGREAT Layer $\times n$
Figure 6: A GREAT layer with sublayers and normalizations
Refer to caption GREAT Input Refer to caption GREAT Edge Encodings Refer to caption Transformed Node Encodings Refer to caption Decoder Solution Generation
Figure 7: A visualization of our GREAT-based RL framework
IV Experiments: Learning to Solve Non-Euclidean Routing Problems
IV-A GREAT-based RL Framework
For our experiments, we built end-to-end RL frameworks for the TSP, CVRP and OP. Our framework follows the encoder-decoder approach (where GREAT serves as the encoder and the multi-pointer network adapted from [8] as the decoder). We train our models to incrementally build solution tours by adding one node at a time to a partial solution.
We point out that we can apply GREAT (which is purely edge-based) to extensions of the TSP such as the CVRP or OP even though capacities and prizes are node-level features. This is because we can easily transform them into edge features. Consider CVRP where a node $j$ has a demand $c_{j}$. We can simply add demand $c_{j}$ to all edges $e_{i,j}$ in the graph. This is because we know that if we have an edge $e_{i,j}$ in the tour, we will visit node $j$ in the next step and therefore need a free capacity in our vehicle big enough to serve the demand of node $j$ which is $c_{j}$. An analogous extension works for OP.
IV-A1 Intuition of the Framework
We visualize the framework in Figure 7 and also show how GREAT potentially learns and passes on information. For the sake of simplicity, we illustrate the idea of the framework for an Euclidean TSP instance. Non-Euclidean instances (and other routing problems) can be processed in the same way. Given an input graph of edge features (e.g., the distances between TSP nodes), GREAT learns edge encodings that reflect the importance of individual edges (i.e., how promising they are to be part of the TSP solution). In the visualization, we assign darker colors to the encodings of such promising edges. Next, the edge encodings are transformed into node encodings by aggregation. This is simply achieved by using a NB GREAT layer that returns the (otherwise temporary) intermediate internal node encodings (note that only the very last GREAT layer returns these node embeddings, otherwise, only edge encodings are passed on to the next layer). Since NF GREAT does not have temporary node embeddings, we use a final NB GREAT layer in our RL framework even when the other GREAT layers are node-free. During this transformation, we assume that the edge information gets passed on and node encodings reflect which nodes are connected by important edges. Our assumption that nodes connected by important edges have similar encodings is backed up by the heatmap visualization in Figure 8. The visualization shows Euclidean distances and cosine similarities between the vector encodings for the nodes in the TSP instance that are returned by the GREAT encoder to be passed on to the decoder. Red frames around a tile in the heatmaps signal that there is an edge between the two nodes in the optimal TSP solution. We can see that for the cosine similarity, the red frames are mostly around tiles with high cosine similarity. Analogously, for the Euclidean distance heatmaps, we can see that the red frames are mostly around tiles with low distances. For generating these heatmaps, we use a node-based GREAT encoder for Euclidean TSP with 100 nodes trained for our experiments. Even though the model is trained on TSP instances of size 100, we can see that the heatmaps indicate similar patterns for TSP instances of size 50 and 30. In Figure 7, we assign similar colors to nodes that are connected by edges deemed promising by GREAT to visualize the effect shown in the heatmaps. In the last step of our encoder-decoder framework, we hypothesize that the decoder can construct solutions from these embeddings by iteratively selecting similar node encodings.
Although it would have been possible to create an edge-based decoder instead of transforming the edge-features produced by the GREAT encoder into node features, converting the edge encodings into node encodings allows us to freely plug in GREAT into existing training pipelines. Further, it reduces the memory requirement from $\mathcal{O}(n^{2})$ edges to $\mathcal{O}(n)$ nodes during decoding.
IV-A2 RL Environment
The rewards for a solution in the environments of the RL framework correspond to the negative traveled distance for TSP and CVRP (ensuring we minimize distance) and to the collected prizes in OP. The rewards of all routing problems are used to formulate losses in a POMO-based setting [7], using multiple rollouts as a robust baseline.
To ensure valid solutions to all considered routing problems, we use masking operations. This means we prevent the model to visit nodes multiple times (TSP, OP and CVRP with exception of the depot), to respect the vehicle capacity (CVRP) and the maximum travel distance (OP). A special case here is the OP in combination with the earlier mentioned XASY distribution. There, it is not simply possible to mask invalid moves that lead to a violation of the maximum travel distance of the OP instance. Since the triangle inequality does not hold for XASY, we cannot rely on simple additive checks to ensure that returning to the depot is still possible. Consider a partial OP solution where we would like to visit a further customer $j$ starting from our current position at node $i$. Even if the direct distances from $i$ to $j$ and from $j$ to the depot appears infeasible (i.e., the sum of these distances would violate the still available distance budget), the actual shortest return path might respect the constraint. Consequently, we would have to compute the true shortest return path to the depot from each candidate node $j$ in order to check feasibility. This would require solving a series of shortest path problems during the decoding process (since after each selection step the shortest path might change as we cannot visit nodes multiple times), which would result in an enormous computational overhead. Therefore, for OP on the XASY distribution, we do not mask moves that could potentially violate maximum travel distance and we require the model to learn valid moves itself. We do this by setting the reward to 0 in case of a violation, incentivizing the model to not take such moves since the goal is to maximize reward.
(a) Cosine similarity TSP30
(b) Cosine similarity TSP50
(c) Cosine similarity TSP100
(d) Euclidean distance TSP30
(e) Euclidean distance TSP50
(f) Euclidean distance TSP100
Figure 8: Vector similarities for node encodings returned by the GREAT encoder in the RL framework for TSP
IV-A3 Experimental Setup
While many other studies focus on developing frameworks to solve larger and larger Euclidean routing problem instances, we mostly focus on instances of “only” $100$ nodes in much more challenging distributions in our experiments. In detail, we train models for TSP, CVRP and OP on the following “distance”-distributions:
  1. 1.
    Euclidean (EUC) distribution where the coordinates are distributed uniformly at random in the unit square.
  2. 2.
    Asymmetric distribution with triangle inequality (TMAT) as used in [9].
  3. 3.
    Extremely asymmetric TSP (XASY) where all pairwise distances are sampled uniformly at random from the interval (0,1).
  4. 4.
    A mixed distribution which consists of $\frac{1}{3}$ EUC, TMAT and XASY, respectively, to obtain trained models applicable to all distributions.
The exact setting for this experiment is the following. For each distribution, we train a GREAT NF and a GREAT NB model with hidden dimension 128. All networks use 5 hidden layers and 8 attention heads. We consider these sizes a reasonable trade-off between expressiveness and resource demands, since larger networks are more powerful but more costly to train and run. We train for for 400 epochs and there are 25,000 instances in each epoch. Every 10 epochs, we change the dataset to a fresh set of 25,000 instances (meaning $400\times 25,000=10,000,000$ instance presentations in total). We used the ADAM optimizer [40] with a learning rate of $0.0001$ to update the network parameters (we found this learning rate to empirically work best in our setting). After each epoch, we evaluate the model on a distinct validation dataset of $1,000$ instances and save the model with the best validation loss during the 400 epochs of training. Furthermore, while training, the distances of all instances in the current data batch are multiplied by a factor in the range (0.5, 1.5) to ensure learning from a more robust data distribution. This allows us to augment the dataset at inference by multiplying the original distances with several factors and solving the resulting scaled instances multiple times. A similar idea was followed in other studies such as [7]. We note that our form of data augmentation by scaling makes the model learn “out of distribution”, much more than the augmentation method used in e.g., [7] which works by rotating Euclidean coordinates. This offers the advantage of potentially increasing the generalization performance of our model as it has encountered distance matrices of very different scales during training. However, this comes at the cost of leading to lower performance gains through augmentation at inference in the original distribution (since the “focus” of the model will by on instances scaled with a factor close to 1, i.e., the original, unscaled instances) compared to the augmentation of [7].
The overall framework to construct solutions, as well as the decoder to decode the encodings provided by GREAT and the loss formulation, are adapted from [8].
We also retrain MatNet [9] and Pointerformer [8] (EUC only) with the exact same hyperparameters and train dataset as GREAT, to create fair benchmarks and comparisons.
IV-B Curriculum Learning
While the main experimental analysis of our work focuses on creating competitive models for TSP, CVRP and OP with 100 nodes on challenging, asymmetric distributions, we also explore tackling larger TSP instances (up to $1000$ nodes) by using a few-shot CL set-up. This means we fine-tune the trained models of TSP for 100 nodes (TSP100) by retraining for one epoch each on the following instance sizes (10% increment steps from the previous instance size and additionally 200 + 500): [110, 121, 133, 146, 161, 177, 194, 200, 214, 235, 259, 285, 313, 345, 379, 417, 459, 500]. For each instance size, we create a dataset of $2,000$ instances. For instance size 200 and 500 we trained for not only one but five epochs (using the same dataset of $2,000$ in each epoch) and the fine-tuned model was saved afterwards. We note that $2,000$ instance presentations is $0.02\%$ of the $10,000,000$ instance presentations that were used to train the “base models” for TSP100, making the fine-tuning “few-shot”. The datasets were used to train the models in the same end-to-end RL framework used for the experiments with 100 nodes. Since the memory requirement and runtime of such end-to-end RL frameworks is quadratic in the number of nodes, it is essential that our model can be fine-tuned using only such small datasets.
While fine-tuning, we kept the POMO size (i.e., the number of samples trajectories, compare [7] for details) to 100 (due to memory constraints), while setting it to the actual instance size during evaluation. We set the batch size to 16 for instance size <250<250, to 8 for instance size <350<350 and 4 otherwise to respect the available VRAM of our used GPU (a single NVIDIA A40 with 48GB of VRAM).
IV-C Training times & Parameters
In the following, we report the training times for training the CVRP models with 100 nodes (other problems might have slightly different times) in Table I. We trained our models for 400 epochs. The total times also includes validating the models after each epoch (on 1000 instances using x8/x9 augmentation) to be able to save the best-performing model after any epoch as our final model and the times for dataset generation (a new train dataset is used every 10 epochs). Therefore, the total time equals Epoch400+Val400+40DataGenEpoch*400+Val*400+40*DataGen where DataGen is approximately 6 minutes (independent of the used model). All training and inference of neural models has been performed on NVIDIA A40 GPUs with 48GB of VRAM. Additionally to training times, we also provide the number of parameters of each model in Table I.
TABLE I: Training times of the considered architectures
Training Times Epoch Val Total Params
GREAT NB $\simeq$ 7.5min $\simeq$ 60s $\simeq$ 60h 1.42M
GREAT NF $\simeq$ 6.9min $\simeq$ 50s $\simeq$ 55h 1.29M
Pointerformer $\simeq$ 2.1min $\simeq$ 12s $\simeq$ 19h 1.09M
MatNet $\simeq$ 3.4min $\simeq$ 22s $\simeq$ 29h 1.42M
In addition to benchmarking GREAT against the retrained MatNet and Pointerformer networks, we also provide several non-learning-based baselines in our experiments. All such baselines were executed on an Apple M4 10-core CPU (at the time of the experiments among the fastest in consumer devices [41]). However, in case of the EA4OP heuristic for the OP, only 1 instance could be solved at a time due the the program running in a docker environment.
TABLE II: Benchmarking GREAT against neural methods and traditional baselines
EUC100 TMAT100 XASY100
opt. gap time opt. gap time opt. gap time
TSP Baseline - Gurobi - - - - - -
LKH 0.0% 442s 0.0% 936s 0.01% 1006s
Nearest Neighbor 24.86% 1.3s 36.04% 1.2s 185.26% 1.2s
Nearest Insertion 21.8% 27s 31.8% 28s 301.65% 28s
Farthest Insertion 7.66% 32s 23.92% 33s 310.98% 33s
BQ-NCO $\textbf{0.01\%}^{*}$ $\text{32m}^{*}$ $\textbf{0.96\%}^{*}$ $\text{19m}^{*}$ UNK UNK
Pointerformer x1 0.96% 18s NA NA NA NA
Pointerformer x8 0.41% 131s NA NA NA NA
MatNet x1 3.28% 29s 7.54% 27s 82.26% 27s
MatNet x9 2.09% 233s 5.28% 229s 62.57% 231s
MatNet x33 1.69% 814s 4.4% 812s 54.55% 808s
GREAT NF x1 1.32% 61s 3.73% 65s 28.54% 64s
GREAT NF x9 0.9% 525s 2.75% 579s 21.15% 568s
GREAT NF x33 0.85% 1952s 2.59% 2071s 20.06% 2075s
GREAT NB x1 1.43% 47s 1.53% 69s 21.93% 70s
GREAT NB x9 1.0% 404s 1.07% 614s 12.6% 610s
GREAT NB x33 0.94% 1446s 1.03% 2235s 10.72% 2206s
GREAT MIX NF x1 4.45% 64s 5.77% 64s 45.78% 64s
GREAT MIX NF x9 3.5% 563s 4.63% 554s 33.32% 552s
GREAT MIX NF x33 3.2% 2017s 4.39% 2014s 29.8% 2015s
GREAT MIX NB x1 2.53% 70s 4.99% 70s 45.97% 70s
GREAT MIX NB x9 1.9% 612s 3.85% 608s 33.19% 609s
GREAT MIX NB x33 1.84% 2212s 3.7% 2211s 30.74% 2207s
Baseline - HGS - 544.2m - 1316.6m - 2163.4m
CVRP Pointerformer x1 4.6% 17s NA NA NA NA
Pointerformer x8 3.29% 115s NA NA NA NA
MatNet x1 4.79% 27s 9.41% 27s 32.54% 27s
MatNet x9 3.6% 205s 7.55% 208s 24.23% 209s
MatNet x33 3.21% 725s 6.72% 727s 20.37% 738s
GREAT NF x1 3.88% 64s 5.49% 65s 12.1% 64s
GREAT NF x9 3.13% 540s 4.4% 551s 6.29% 541s
GREAT NF x33 3.01% 1958s 4.2% 1983s 5.31% 1963s
GREAT NB x1 3.06% 71s 4.66% 70s 8.19% 70s
GREAT NB x9 2.47% 601s 3.72% 594s 2.23% 595s
GREAT NB x33 2.39% 2184s 3.6% 2172s 0.76% 2171s
GREAT MIX NF x1 4.12% 64s 6.37% 64s 14.83% 64s
GREAT MIX NF x9 3.38% 543s 5.11% 536s 8.04% 542s
GREAT MIX NF x33 3.15% 1961s 4.8% 1956s 5.42% 1962s
GREAT MIX NB x1 3.98% 70s 5.49% 71s 11.88% 71s
GREAT MIX NB x9 3.18% 594s 4.35% 598s 4.91% 604s
GREAT MIX NB x33 3.03% 2148s 4.15% 2183s 3.18% 2199s
Baseline - Greedy - 7.9s - 6.9s - 340s
OP EA4OP -23.64% 5609s NA NA NA NA
Pointerformer x1 -19.61% 12s NA NA NA NA
Pointerformer x8 -21.31% 74s NA NA NA NA
MatNet x1 -15.54% 21s -24.13% 21s 54.98% 17s
MatNet x9 -18.08% 161s -25.51% 166s 28.76% 137s
MatNet x33 -18.99% 551s -26.01% 583s 20.15% 471s
GREAT NF x1 -20.92% 59s -27.45% 60s -18.1% 56s
GREAT NF x9 -21.82% 498s -28.24% 510s -21.34% 479s
GREAT NF x33 -22.03% 1806s -28.39% 1840s -21.5% 1738s
GREAT NB x1 -21.65% 65s -28.08% 66s -20.69% 62s
GREAT NB x9 -22.38% 552s -28.71% 559s -24.18% 533s
GREAT NB x33 -22.53% 1999s -28.87% 2011s -24.46% 1948s
GREAT MIX NF x1 -16.97% 58s -24.43% 60s -15.84% 55s
GREAT MIX NF x9 -18.59% 493s -26.07% 498s -21.06% 478s
GREAT MIX NF x33 -19.09% 1774s -26.38% 1806s -21.78% 1742s
GREAT MIX NB x1 -18.22% 65s -25.5% 66s -19.34% 62s
GREAT MIX NB x9 -19.56% 548s -27.02% 556s -23.7% 537s
GREAT MIX NB x33 -19.96% 1995s -27.35% 2032s -24.08% 1928s
* values from original paper, model was not reevaluated by us
TABLE III: Generalization results on test datasets with 30 instances from [34]
TMAT150 TMAT250 TMAT1000
opt. gap time opt. gap time opt. gap time
LKH - 6s - 15s - 231s
BQ-NCO bs16 UNK UNK UNK UNK $8.26\%^{*}$ $7\text{m}^{*}$
GLOP $19.28\%^{\text{\textdagger}}_{\diamond}$ $8.2s^{\text{\textdagger}}$ $29.67\%^{\text{\textdagger}}_{\diamond}$ $9.3s^{\text{\textdagger}}$ $44.27\%^{\text{\textdagger}}_{\diamond}$ $15s^{\text{\textdagger}}$
GREAT NF ZS x1 4.47% 3s 11.98% 6s 85.54% 54s
GREAT NF ZS x9 3.42% 19s 7.81% 17s 52.07% 455s
GREAT NB ZS x1 1.64% 3s 2.23% 6s 4.17% 56s
GREAT NB ZS x9 1.28% 11s 1.8% 17s 2.76% 470s
GREAT NF FT200 x1 4.3% 3s 4.91 % 6s 19.05% 54s
GREAT NF FT200 x9 3.28% 8s 4.18% 17s 15.42% 457s
GREAT NB FT200 x1 2.12% 3s 2.6% 6s 1.8% 56s
GREAT NB FT200 x9 1.54% 8s 1.9% 18s 1.45% 472s
values from the original paper (different dataset of 128 instances)
$^{\text{\textdagger}}$ values from the original paper (same dataset as GREAT)
$\diamond$ original paper reports absolute values which we transformed into opt. gaps
TABLE IV: Generalization results on test datasets with 128 instances
EUC TMAT XASY
opt. gap time opt. gap time opt. gap time
TSP200 Baseline - LKH - 23s - 40s - 36s
Nearest Neighbor 25.5% 0.1s 36.71 % 0.1s 224.93 % 0.1s
Nearest Insertion 23.39% 3.8s 37.95 % 3.9s 465.76% 4.1s
Farthest Insertion 9.04% 2.9s 29.41 % 2.9s 481.4% 3.0s
BQ-NCO bs16* 0.09%* 3min* 1.41%* 1min* UNK UNK
GREAT NF ZS x1 15.78% 12s 6.97% 12s 209.3 % 12s
GREAT NF ZS x9 13.69% 41s 5.42% 44s 148.18% 43s
GREAT NB ZS x1 5.27% 11s 2.27% 13s 53.28 % 13s
GREAT NB ZS x9 4.54% 37s 1.77% 55s 37.47 % 53s
GREAT NF FT200 x1 3.29% 13s 5.08% 12s 56.0 % 12s
GREAT NF FT200 x9 2.66% 37s 4.18% 38s 45.22 % 38s
GREAT NB FT200 x1 3.47% 11s 2.48% 13s 51.43 % 13s
GREAT NB FT200 x9 2.86% 29s 1.92% 42s 38.38% 42s
GREAT NF FT500 x1 105.86% 37s 7.28% 12s 142.91% 12s
GREAT NF FT500 x9 92.58% 13s 6.46% 38s 103.47% 38s
GREAT NB FT500 x1 5.94 % 29s 29.08% 13s 134.77 % 13s
GREAT NB FT500 x9 4.67 % 10s 28.54% 42s 127.68 % 42s
TSP500 Baseline - LKH - 272s - 373s - 367s
Nearest Neighbor 25.78% 0.4s 37.27 % 0.4s 280.26 % 0.5s
Nearest Insertion 25.12% 41s 46.77% 51s 801.73% 49s
Farthest Insertion 10.66% 50s 36.61% 65s 815.74 % 77s
BQ-NCO bs16* 0.55%* 15min* 2.43%* 3min* UNK UNK
GREAT NF ZS x1 49.81% 49s 56.46% 46s 369.13 % 46s
GREAT NF ZS x9 39.51% 288s 31.42% 301s 309.12 % 297s
GREAT NB ZS x1 16.75 % 40s 3.66% 49s 307.52 % 49s
GREAT NB ZS x9 15.89% 247s 3.14% 315s 246.44 % 318s
GREAT NF FT200 x1 36.13% 46s 7.72 % 45s 470.31 % 45s
GREAT NF FT200 x9 33.13% 286s 6.81% 291s 367.82 % 293s
GREAT NB FT200 x1 10.96 % 39s 3.42% 49s 136.48 % 46s
GREAT NB FT200 x9 10.16 % 245s 3.01% 317s 118.01% 311s
GREAT NF FT500 x1 7.52% 44s 7.06% 45s 128.75% 46s
GREAT NF FT500 x9 6.96% 283s 6.5% 291s 105.23% 314s
GREAT NB FT500 x1 8.04 % 42s 32.99% 46s 179.01 % 47s
GREAT NB FT500 x9 7.11 % 259s 31.98% 310s 167.71 % 310s
values from original paper tested on a different dataset of 128 instances
TABLE V: Generalization results on real world instances from TSPLIB, CVRPLIB and OPLIB (1 to 100 nodes)
Dist. Mean Q10 Q25 Median Q75 Q90
TSPLIB $n=12$ Baseline - LKH - - - - - - -
GREAT NB x9 EUC 2.38% 0.66% 1.32% 2.16% 3.44% 3.65%
GREAT NF x9 EUC 20.45% 0.72% 1.9% 3.56% 7.17% 57.38%
GREAT NB x9 MIX 2.25% 1.45% 1.85% 2.28% 2.59% 2.82%
GREAT NF x9 MIX 5.6% 1.61% 2.69% 4.64% 7.76% 10.68%
MatNet x9 EUC 2.82% 1.08% 1.43% 3.08% 3.55% 3.97%
Pointerformer x8 EUC 1.36% 0.04% 0.18% 1.19% 2.33% 2.87%
ATSPLIB $n=14$ Baseline - LKH - - - - - - -
GREAT NB x9 TMAT 3.81% 0.46% 0.97% 2.65% 6.0% 8.85%
GREAT NF x9 TMAT 6.03% 2.1% 3.47% 6.06% 8.56% 10.4%
GREAT NB x9 XASY 45.88% 9.33% 14.74% 18.27% 41.47% 80.41%
GREAT NF x9 XASY 27.12% 6.26% 18.48% 27.32% 37.51% 39.37%
GREAT NB x9 MIX 17.65% 5.22% 8.46% 15.79% 24.29% 29.97%
GREAT NF x9 MIX 19.58% 4.65% 5.53% 7.44% 9.06% 16.09%
MatNet x9 TMAT 16.93% 8.89% 12.29% 18.94% 20.1% 25.32%
MatNet x9 XASY 14.46% 5.09% 9.62% 15.04% 20.43% 22.06%
CVRPLIB $n=92$ Baseline - HGS - - - - - - -
GREAT NB x9 EUC 3.82% 1.44% 2.05% 2.93% 4.63% 7.52%
GREAT NF x9 EUC 50.49% 4.62% 7.37% 30.24% 92.81% 120.92%
GREAT NB x9 MIX 3.92% 1.46% 2.23% 3.24% 4.46% 7.49%
GREAT NF x9 MIX 7.93% 2.55% 3.32% 4.83% 8.02% 15.21%
MatNet x9 EUC 7.62% 2.04% 3.03% 4.01% 8.09% 16.57%
Pointerformer x8 EUC 4.09% 1.33% 2.2% 3.21% 4.47% 7.01%
OPLIB $n=48$ Baseline - EA4OP - - - - - - -
GREAT NB x9 EUC 8.68% 1.83% 2.84% 5.18% 6.46% 8.29%
GREAT NF x9 EUC 11.66% 3.59% 4.75% 8.37% 10.58% 28.11%
GREAT NB x9 MIX 9.12% 4.19% 5.64% 8.92% 11.14% 13.0%
GREAT NF x9 MIX 12.1% 6.31% 8.91% 11.04% 13.42% 18.64%
MatNet x9 EUC 14.9% 7.21% 10.76% 13.26% 17.37% 24.36%
Pointerformer x8 EUC 8.19% 2.0% 4.81% 7.5% 9.87% 13.49%
TABLE VI: Generalization results on real world instances from TSPLIB, CVRPLIB and OPLIB (101 to 200 nodes)
Dist. Mean Q10 Q25 Median Q75 Q90
TSPLIB $n=17$ Baseline - LKH - - - - - - -
GREAT NB x9 EUC 5.97% 1.68% 2.2% 3.94% 6.89% 10.76%
GREAT NF x9 EUC 9.27% 1.55% 3.44% 7.07% 14.26% 22.3%
GREAT NB x9 MIX 6.28% 2.42% 3.11% 4.35% 7.03% 8.56%
GREAT NF x9 MIX 14.26% 3.81% 5.81% 13.87% 20.91% 23.64%
Pointerformer x8 EUC 3.9% 0.93% 1.63% 2.18% 5.3% 6.88%
CVRPLIB $n=49$ Baseline - HGS - - - - - - -
GREAT NB x9 EUC 7.97% 2.38% 2.81% 5.23% 6.72% 10.49%
GREAT NF x9 EUC 9.38% 3.64% 5.87% 8.43% 11.9% 15.87%
GREAT NB x9 MIX 14.94% 3.57% 4.68% 6.23% 9.8% 12.01%
GREAT NF x9 MIX 9.36% 3.55% 4.9% 7.41% 12.92% 17.5%
Pointerformer x8 EUC 8.61% 3.77% 5.14% 6.25% 12.4% 15.46%
OPLIB $n=60$ Baseline - EA4OP - - - - - - -
GREAT NB x9 EUC 18.54% 2.64% 4.81% 6.11% 9.91% 17.94%
GREAT NF x9 EUC 16.16% 5.61% 7.57% 11.81% 17.94% 24.86%
GREAT NB x9 MIX 14.39% 7.17% 8.44% 11.16% 15.34% 30.06%
GREAT NF x9 MIX 19.36% 7.46% 11.59% 16.84% 24.07% 30.99%
Pointerformer x8 EUC 12.93% 5.95% 7.49% 10.52% 13.02% 19.1%
IV-D Results
IV-D1 TSP, CVRP and OP with 100 nodes
We provide an overview of the performance of our models and their inference times in Table II. The results show the average performance and total runtime on $10,000$ test instances. We use Gurobi [42] as our baseline for TSP, HGS [17] for CVRP and a custom greedy heuristic that trades off the collectible prize with the distance for OP (compare Appendix Acknowledgments). We use this greedy baseline for OP due to the absence of other good heuristics that can be applied to all variants of OP (i.e., including asymmetric OP) out of the box. Only for EUC OP, we can also use and apply the the EA4OP heuristic [18] and provide these values as an additional comparison. For TSP, we also report results for LKH [15] and some simple heuristics, as well as the very good results achieved by the neural approach BQ-NCO [35]. However, we did not retrain BQ-NCO (since the overall pipeline and framework is completely different from ours) and simply report the values provided in the original paper.
For EUC TSP, BQ-NCO is the best neural approach, followed by Pointerformer, GREAT NF, GREAT NB and MatNet. We note that GREAT NB is faster here than GREAT NF (and the GREAT NB in other experimental settings) since we have a completely symmetric graph and the attention score computation simplifies to a single expression (compare Section III-B). We point out that the results for MatNet and Pointerformer deviate from the original studies since we retrained them on a smaller dataset (same as the GREAT training dataset) and with adjusted hyperparameters (same as for GREAT) to make the comparison fair. On TMAT and XASY TSP, GREAT is the second and best performing model, respectively, achieving gaps of $1.03\%$ and $10.72\%$ with a $\times 33$ augmentation. We note that GREAT performs considerably better than MatNet on these distributions while Pointerformer is not applicable at all due to its limitation to Euclidean inputs. We point out, however, that MatNet (and also Pointerformer) are generally faster than GREAT. We want to highlight the large remaining gap on XASY, showing the complexity of the learning task on this distribution.
On CVRP, GREAT generally achieves the best results. However, for EUC CVRP there are other promising neural models achieving better results, e.g. [35] or also [27]. These are, however, applied only to EUC CVRP and to the best of our knowledge, we are the first ones to learn to solve CVRP with the TMAT and XASY distribution. We note that the runtimes of GREAT are much faster than HGS while achieving relatively small gaps, especially for XASY. This is surprising giving the rather large gaps for XASY TSP. We hypothesize that this is due to the customer demands and vehicle capacity, resulting in several subtours that impose more structure on the potential solution.
GREAT is again our best performing model, but we note that on EUC OP [35] achieves better performance than us (as well as the non-learning based EA4OP heuristic). Again, to the best of our knowledge, this work is the first to learn to solve OP on the TMAT and XASY distributions. Notably, this is significant because, as far as we are aware, no advanced heuristics currently exist for accurately solving the asymmetric OP.
Overall, for problems with 100 nodes, we summarize that GREAT is generally among the best performing models in our set-up. Notably, for EUC, there are other neural models available that produce slightly better results, however, they are typically limited to Euclidean settings and, to the best of our knowledge, we are the first ones to learn to solve asymmetric CVRP and OP. Among the two different GREAT versions, GREAT NB seems to generally perform better, with the exception of EUC TSP. Further, for most problems (with the exception of TSP XASY), GREAT MIX performs almost as well as the pure GREAT models, showing that it is possible for a single GREAT model to learn all distributions at the same time. We hypothesize increasing the training dataset size could close these gaps since a GREAT MIX model encounters only $\frac{1}{3}$ of the data of a particular distribution during training compared to a “specialized” model. We hypothesize that GREAT is unable to unfold its full potential (compared to existing node-level models which, as we acknowledged, often perform slightly better on EUC) due to our scaling-based augmentation method which is disadvantageous compared to Euclidean coordinate flipping when focusing on a single distribution.
IV-D2 Generalization to Large Instances and Curriculum Learning
We now investigate the generalization performance of GREAT to larger instances. For this, we report two different types of results: firstly, the generalization performance of the un-fine-tuned original model trained for TSP with 100 nodes (TSP100) in a zero-shot (ZS) generalization and, secondly, the performance of models that were fine-tuned in a few-shot curriculum learning framework for TSP with 200 cities and 500. We denote the latter models with FT200 and FT500, respectively. We report the results when evaluating on 128 test instances in Table IV. We use LKH as our baseline and further provide the results of three basic heuristics and [35] for comparison.
We note that, in contrast to the main experiments on problems with 100 nodes, in this setting GREAT NF often achieves the best performance among all GREAT models (for TSP500). We observe that GREAT NF typically seems to perform best on the actual fine-tuned distribution. However, GREAT NB seems to have a very strong generalization performance (especially on the asymmetric distributions) towards larger instances than it was trained on. This can be seen by the ZS and FT200 variants achieving better performance than the variants fine-tuned for larger instances. We find this very surprising and attribute it to our scaling-based augmentation methods (which we also apply during training). While this method might be suboptimal in a fixed setting of instances with exactly 100 nodes, it seems to give our GREAT models good generalization capabilities. We hypothesize that a smaller TSP instance can look similar to a subregion of a larger TSP instance with the right scaling, making it easier for the model to generalize.
As an additional comparison, we evaluate GREAT on the same large TMAT test instances used in GLOP [34] in Table III (with 150, 250 and 1000 nodes). We can see that GREAT NB achieves convincing performance in the zero-shot setting and when fine-tuned on instances of size 200, outperforming GLOP. This is surprising and shows strong generalization performance of GREAT, keeping in mind that GLOP is a divide-and-conquer-based approach designed to work on larger instances than the one it was trained on.
IV-D3 TSPLIB, CVRPLIB and OPLIB results
In addition to the experiments conducted on synthetic data, we also test our trained models on real world instances of the TSPLIB [33], CVRPLIB [43] and OPLIB [^2] https://github.com/bcamath-ds/OPLib/. We note that we only consider instances with at most $200$ nodes (for MatNet, the limit is $100$ nodes since we use a hidden dimension of $128$ in all our models, which means our MatNet can only produce initial one-hot encodings for instances with up to $128$ nodes). For OPLIB and CVRPLIB, we only consider Euclidean instances. For TSPLIB, we consider Euclidean and asymmetric instances. The results for instances with $1$ to $100$ nodes can be found in Table V, the results for instances with $101$ to $200$ nodes in Table VI. The tables specify the amount of instances that are considered for a given library, e.g., there are $n=12$ Euclidean TSP instances in TSPLIB with at most $100$ nodes. Since the instances in the various real-world libraries can differ significantly, we report not only the mean optimality gap (as we do for synthetic data), but also the median, as well as the 10th (q10), 25th (q25), 75th (q75), and 90th (q90) percentiles of the optimality gaps. Note that optimality gaps are computed w.r.t. LKH for TSP, HGS for CVRP and EA4OP for OP. Additionally, the tables also have a column Dist. which specifies the distribution of the training dataset that was used to train the evaluated models.
From a performance perspective, Pointerformer achieves the best results on Euclidean TSPLIB for all considered instance sizes. This is in line with the results on Euclidean synthetic data. For asymmetric TSP instances, the node-based GREAT trained on the TMAT distribution achieves the best results. Note that for asymmetric TSP, there is only a single instance with more than 100 nodes and we cannot benchmark against MatNet (due to the one-hot encoding limitation) nor Pointerformer (limited to Euclidean coordinates) so we omit this experiment. For CVRPLIB instances with 100 nodes and less, node-based GREAT trained on the EUC distribution leads for the mean, q25 and median scores, but Pointerformer achieves the better 10th and 90th percentile results. For instances, with 101 to 200 nodes, however, GREAT NB trained on EUC leads for all statistics. On OPLIB instances (for all considered sizes), Euclidean node-based GREAT leads the table for all statistics but the mean, where Pointerformer performs best. This is most likely due to a few outlier instances where GREAT performs exceptionally poorly, which disproportionately affect the mean. This is supported by the fact that the mean of GREAT NB (EUC) is worse than its 90th percentile score. Overall, we summarize that GREAT shows strong generalization performance on the considered real world test instances, with Pointerformer outperforming GREAT only on Euclidean TSP data, similarly to the tests on synthetic data.
V Conclusion
In this work, we introduce GREAT, a novel GNN-related neural architecture for edge-level graph problems. Compared to existing architectures, GREAT offers the advantage of directly operating on the edge features. In the context of routing problems, operating on edge features such as distances overcomes the limitation of previous Transformer and GNN-based models that operate on node coordinates which essentially limits these architectures to Euclidean problems. This limitation is rather disadvantageous in real-life settings, however, as distances (and especially other characteristics like time and energy consumption) are often asymmetric due to topography (e.g., elevation) or traffic congestion.
To validate GREAT, we develop a GREAT-based RL framework to solve vehicle routing problems like TSP, CVRP and OP. GREAT achieves competitive performance on several distributions of these routing problems (EUC, TMAT and XASY) and, to the best of our knowledge, is the first neural model to learn asymmetric versions of CVRP and OP altogether. In particular, for instances with 100 nodes, we achieve gaps of $3.6\%$ for CVRP on TMAT, $0.76\%$ for CVRP on XASY, $28.87\%$ for OP on TMAT and $24.46\%$ for OP on XASY compared to our baselines. GREAT also shows promising generalization performance on larger TSP instances in zero-shot generalization without any fine-tuning and, even better performance after a few-shot CL pipeline where the instance sizes are incrementally increased. Further, we confirm that GREAT, while trained on synthetic data, achieves promising performance in real world symmetric and asymmetric TSP instances, as well as CVRP and OP instances.
In future extensions, we are interested in investigating how GREAT can be used as a plug-in method for divide-and-conquer-based approaches like GLOP [34] that make it possible to solve much larger problem instances compared to our current end-to-end RL framework. We further believe that GREAT could be useful in edge-classification and edge-regression tasks (e.g., in the setting of [3]) and, possibly, beyond routing problems.
Acknowledgments
This work was performed as a part of the research project “LEAR: Robust LEArning methods for electric vehicle Route selection” funded by the Swedish Electromobility Centre (SEC). The computations were enabled by resources provided by the National Academic Infrastructure for Supercomputing in Sweden (NAISS) at Chalmers e-Commons partially funded by the Swedish Research Council through grant agreement no. 2022-06725.
[Greedy Heuristic for Orienteering Problem] In our experiments, we use a greedy heuristic as a baseline for solving the OP. This method prioritizes visiting customers that offer a favorable trade-off between prize and distance, while respecting the specified maximum travel distance. At each step, starting from the current node $i$, the algorithm selects the next unvisited node $j$ that maximizes the ratio
$\frac{\text{prize}(j)}{\text{distance}(i,j)},$
subject to the constraint that traveling to $j$ and then returning to the depot does not exceed the remaining travel budget. This means that we mask all moves where
$\text{distance}(i,j)+\text{distance}(j,\text{depot})$
exceeds the remaining travel distance.
While this approach is straightforward on Euclidean instances, applying it to instances of the asymmetric XASY distribution is significantly more challenging. Since the triangle inequality does not hold for XASY, we cannot rely on simple additive checks to ensure feasibility. That is, even if the direct distances from $i$ to $j$ and from $j$ to the depot appears infeasible, the actual shortest return path might respect the travel budget. Consequently, we must compute the true shortest return path to the depot from each candidate node $j$ in order to check feasibility. This requires solving a series of shortest path problems during the greedy selection process (since after each selection step the shortest path might change as we cannot visit nodes multiple times), which results in an increased runtime on XASY instances.